# Quantum Entanglement Trees: Optimizing Quantized Matrix Quantization via Element Replacement and Residual Clustering

### v3

@@ Abstract
Experimental results demonstrate that QET can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms. Our contributions include the abstraction of the QEM problem, the design of the QET algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@ 1 Introduction
We evaluate QET on multiple datasets. Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
@@
  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate QET algorithm.
@@ 3.3 Optimizations
In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO). RO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE. On the other hand, BPO utilizes batch processing techniques to enhance processing speed. Since the codebook used in the QET algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
@@
Compression Ratio 4.0 8.0 12.0 16.0
Range Ratio (%) 1.6 4.0 9.2 29.7
CB Proportion (%) 87.5 78.1 71.8 62.5
3.3.1 Residual Optimization
Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm. As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range. Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO). The RO algorithm, as depicted in Figure 2, begins with the quantization of the original matrix to produce a quantized matrix. This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix. The residual matrix typically exhibits a significantly reduced data range, making it easier to compress. This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or QET.
@@ 3.3.1 Residual Optimization
3.3.2 Batch Processing Optimization
The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the QET algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
In the original process, elements of the matrix are compared and sorted individually, leading to a sequential, element-by-element approach that can become computationally intensive as the matrix size increases. This method is not optimal for large-scale matrices.
BPO introduces a more advanced technique by processing matrix elements in parallel batches rather than sequentially. This parallelization allows the algorithm to handle multiple comparisons and sorting operations simultaneously, rather than one at a time. By doing so, it reduces the overall computational complexity and accelerates the entire quantization process.
3.3.3 Codebook Quantization Optimization (CQO)
The codebook occupies a significant portion of the space in the QET algorithm. As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space. Therefore, compressing the codebook can significantly reduce the space used by the algorithm. Our optimization mainly employs the RTN method to compress the codebook.
4 Theoretical Guarantees of QET
Theorem 1.
Suppose each element in the matrix is independently sampled from a normal distribution $x_{(i,j)}$ $\sim$ $\mathcal{N}(\mu,\sigma)$. After Quantizated by the PQ algorithm, the MSE is as follows:
$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$
$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
[…]
$C_{ij}\quad\text{residual matrix}$
$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
$\text{MSE}=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (1)
In this section, this paper will theoretically prove that the MSE of our HERA algorithm is lower than that of the PQ algorithm.
Suppose each element in the matrix is independently sampled from a normal distribution $\mathcal{N}(\mu,\sigma)$. After compression by the PQ algorithm, the process for calculating MSE is as follows:
$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
[…]
$C_{ij}\quad\text{residual matrix}$
$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
$MSE=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
$=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (2)
For the HERA algorithm proposed in this paper, the MSE obtained using the HERA algorithm on a matrix with elements following a $\mathcal{N}(\mu,\sigma^{2})$ distribution is equivalent to the MSE obtained using the PQ algorithm on a matrix with elements following a distribution with variance $0.682\sigma^{2}$. According to formula equation 2 equation 3, the MSE of HERA is smaller than that of the PQ . The following discussion in this paper will explain why this equivalence holds.
For a Gaussian distribution, after applying element pairing in the HERA algorithm, the cumulative distribution function $F_{Y}(y)$ of the elements on the right side of the matrix (the side storing the larger elements) satisfies:
$F_{Y}(y)=P(Y\leq y)=P(\max(X_{1},X_{2})\leq y)=P(X_{1}\leq y)\cdot P(X_{2}\leq y)$
$\text{Each}\ X_{i}\text{ follows a normal distribution}\ N\ (\mu,\sigma^{2})$
Take the derivative of $\text{Take the derivative of }F_{Y}(y):$
The mean and variance of $\text{The mean and variance of }Y\text{ are given by:}$:
[…]
Use numerical integration methods:
$E(Y)=\mu+0.564\sigma$
$Var(Y)=0.682\sigma^{2}$ (3)
Taking symmetry into account, the left side of the matrix also adheres to equation equation 3.
@@ 5.1 Experiment setup
Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset. Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text. The results for the other datasets are provided in the appendix (Section LABEL:sec:appexponaddi). Below, we provide a detailed description of each dataset.
(1) Synthetic normal distribution dataset: This dataset was generated by drawing each element of the matrix from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16. Additionally, one out of every ten thousand elements was replaced with an outlier value, randomly chosen between -100 and 100.
(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al. (2023)).
(3) KV cache dataset: The KV cache dataset is derived from the key-value pairs stored in the cache during inference in KV Quant (Hooper et al. (2024)). The key and value are stored in matrices, respectively.
(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks. GIST consists of the first 100k features extracted from the tiny image set (Torralba et al. (2008)). The dataset is also used in Jegou et al. (2010) and Oliva & Torralba (2001).
(5) Wiki dataset: The Wiki dataset (wik (2023)) contains document embeddings derived from Wikipedia articles. This dataset includes 256-dimensional vectors for a sample of 1 million Wikipedia entries.
Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Core i9-10980XE processor, featuring 18 cores and 36 threads, operating at a base frequency of 3.00 GHz. The server also includes 128GB of 3200MHz DDR4 memory and a 24.8MB L3 cache, providing robust computational capabilities. All algorithms were implemented in Python, using version 3.8.10. For each case, the experiment was repeated 100 times. In each repetition, matrices of the same size and from the same distribution were generated using different seeds.
Metrics: We primarily measure the accuracy and time consumption of the algorithm. We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) as accuracy metrics. QTR (Quantization Time Ratio) and DTR (Dequantization Time Ratio) are used as time consumption metrics. Below, we introduce these metrics in detail.
@@
For time efficiency, we define the following metrics:
$\text{QTR}=\frac{T_{\text{quant}}}{T_{\text{quant, baseline}}},\;\text{DTR}=\frac{T_{\text{dequant}}}{T_{\text{dequant, baseline}}}$
In these equations, $T_{\text{quant}}$ and $T_{\text{dequant}}$ represent the time taken by the proposed algorithm to quantize and dequantize the data, respectively. $T_{\text{quant, baseline}}$ and $T_{\text{dequant, baseline}}$ refer to the corresponding times taken by the baseline algorithm.
@@
Parameter Selection
In this section, we describe the parameter selection process for our HERA algorithm and compare its performance with the Optimized Product Quantization (OPQ) algorithm.
For Residual Optimization (RO), the large space required by RTN makes it infeasible for the algorithm to achieve high compression ratios. Therefore, in our experiments, we primarily use QET as the residual algorithm for RO.
Our QET algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
where:
  • $k_{s}$ is the number of centroids,
  • $d$ is the dimensionality of each sample,
  • $n$ is the number of samples,
  • $m$ is the number of subspaces,
  • bitlength is the bit length used for encoding.
This constraint ensures that the total memory usage remains within the available memory limits. By carefully selecting $k_{s}$, we aim to optimize the performance of the HERA algorithm while adhering to memory constraints.
The Optimized Product Quantization (OPQ) algorithm was used as a benchmark for comparison. The OPQ algorithm was implemented with default parameters, providing a consistent basis for performance evaluation. The OPQ algorithm optimizes product quantization by adjusting the space partitioning and centroid assignments to minimize quantization error.
Accuracy Measurement
Figure 3: MAE on 8 subspaces.
Figure 4: MRE on 8 subspaces.
Figure 5: MSE on 8 subspaces.
Figure 6: MAE on 16 subspaces.
Figure 7: MRE on 16 subspaces.
Figure 8: MSE on 16 subspaces.
Figure 9: MAE on 32 subspaces.
Figure 10: MRE on 32 subspaces.
Figure 11: MSE on 32 subspaces.
Figure 12: MAE on 128 subspaces.
Figure 13: MRE on 128 subspaces.
Figure 14: MSE on 128 subspaces.
Figure 15: MAE on 1 iterations.
Figure 16: MRE on 1 iterations.
Figure 17: MSE on 1 iterations.
Figure 18: MAE on 2 iterations.
Figure 19: MRE on 2 iterations.
Figure 20: MSE on 2 iterations.
Figure 21: MAE on 3 iterations.
Figure 22: MRE on 3 iterations.
Figure 23: MSE on 3 iterations.
Figure 24: MAE on 4 iterations.
Figure 25: MRE on 4 iterations.
Figure 26: MSE on 4 iterations.
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  5-  5, Figures  8-  8, Figures  11-  11, Figures  14-  14 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
From the experimental results, we can derive four conclusions. Firstly, the higher the compression ratio, the less space is consumed, but the accuracy decreases. Secondly, the OPQ algorithm performs worse than others with the same compression rate because it needs to store the transformation matrix. Additionally, the PQ algorithm does not perform as well as our improved algorithm. Finally, the more iterations, the better the performance of our algorithm.
We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1717, Figures  2020, and Figures  2323, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).

### v4

# QET: Enhancing Quantized LLM Parameters and KV cache Compression through Element Substitution and Residual Clustering

@@ Abstract
Experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the current best method on the LLM dataset, K cache, and V cache, respectively. Our contributions include the abstraction of the QEM problem, the design of the QET algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@ 1 Introduction
We evaluate QET on multiple datasets. Experimental results demonstrate that QET can effectively reduce MSE to 5.05%, 13.33%, and 11.89% of the current best method on the LLM dataset, K cache, and V cache, respectively.
@@
  • Proposed two optimizations: We introduced residual quantization to reduce MSE and codebook quantization techniques to optimize memory efficiency.
@@ 3.2.1 QET Dequantization
3.3 Theoretical Guarantees of QET
Theorem 1.
Suppose each element in the matrix is independently sampled with a mean $\mu$ and variance $\sigma^{2}$. After quantization by the PQ algorithm, the Mean Squared Error (MSE) is given by:
$\text{MSE}_{\text{PQ}}=\sigma^{2}\left(1-\frac{1}{n_{k}}\right).$
Proof.
Let $c_{(i,j)}$ denote the cluster centroid matrix, and let $n_{k}$ be the number of vectors in the cluster. Define the residual matrix as $s_{(i,j)}$.
The expectation of the residual matrix is:
$E[s_{(i,j)}]=E[x_{(i,j)}-c_{(i,j)}]=\mu-\mu=0.$
The Mean Squared Error (MSE) can be expressed as the variance of the residual matrix:
$\text{MSE}_{\text{PQ}}=\text{Var}[s_{(i,j)}]=\text{Var}[x_{(i,j)}]+\text{Var}[c_{(i,j)}]-2\cdot\text{Cov}(x_{(i,j)},c_{(i,j)}).$
Given that $c_{(i,j)}$ is the centroid within the cluster:
$\text{MSE}_{\text{PQ}}$ $=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(x_{(i,j)}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}x_{(i,j)}-\mu\right)\right]$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{1}{n_{k}}\left(E\left[(x_{(i,j)}-\mu)^{2}\right]+\sum_{i\neq i^{\prime}}E\left[(x_{(i,j)}-\mu)(x_{i^{\prime}j}-\mu)\right]\right).$
Since $x_{(i,j)}$ and $x_{i^{\prime}j}$ (for $i\neq i^{\prime}$) are independent:
$\text{MSE}_{\text{PQ}}=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}=\left(1-\frac{1}{n_{k}}\right)\sigma^{2}.$ (1)
Theorem 2.
For the QET algorithm without optimization, where matrix elements are independently sampled from a normal distribution $x_{(i,j)}\sim\mathcal{N}(\mu,\sigma^{2})$, the Mean Squared Error (MSE) is:
$\text{MSE}_{\text{QET}}=0.682\times\text{MSE}_{\text{PQ}}.$
Proof.
We denote by $F_{Y}(y)$ the cumulative distribution function (CDF) of the elements on the right side of the matrix (i.e., the larger elements).
$F_{Y}(y)=P(Y\leq y)=P(\max(X_{l},X_{r})\leq y)=P(X_{l}\leq y)\cdot P(X_{r}\leq y)$
Where $y$ represents the elements of the matrix formed by the larger of the adjacent elements in matrix $X$. $X_{l}$ and $X_{r}$ denote the random variables corresponding to the adjacent left and right elements in matrix $X$, respectively. Since $X_{l}$ and $X_{r}$ follow a normal distribution $\mathcal{N}(\mu,\sigma^{2})$, we have:
Take the derivative of $F_{Y}(y)$ with respect to $y$:
The mean and variance of $Y$ are given by:
[…]
Using numerical integration methods, we obtain:
$\mathbb{E}(Y)=\mu+0.564\sigma,\quad\text{Var}(Y)=0.682\sigma^{2}.$
Because QET can be considered a special case of the PQ algorithm, according to Equation 1,
$\text{MSE}_{\text{QET}}=\left(1-\frac{1}{n_{k}}\right)\text{Var}(Y)=0.682\left(1-\frac{1}{n_{k}}\right)\sigma^{2}=0.682\cdot\text{MSE}_{\text{PQ}}.$ (2)
Taking symmetry into account, this conclusion also holds for the smaller matrix on the left. ∎
Theorem 1 demonstrates that the MSE of quantized matrix is proportional to the variance of the distribution. For any distribution, as long as the variance of the matrix elements is reduced, the MSE will also decrease. Theorem 2 shows that when the matrix elements follow a normal distribution, the MSE is reduced to 0.682 times its original value.
@@ 3.4 Optimizations
In this section, we propose three optimizations: Residual Quantization Optimization (RQO) and Codebook Quantization Optimization (CQO). RQO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE. CQO further quantizes the codebook to improve space efficiency.
@@
Compression Ratio 4.0 8.0 12.0
Range Ratio (%) 1.6 4.0 9.2
CB Proportion (%) 87.5 78.1 71.8
3.4.1 Residual Quantization Optimization
Residual Quantization Optimization (RQO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm. As shown in Table 1, our experiments indicate that at a compression ratio [^3] Compression ratio = Space occupied by the matrix before quantizationSpace occupied by the matrix after $\frac{\text{Space occupied by the matrix before quantization}}{\text{Space occupied by the matrix after quantization}}$ of 12, the data range of the quantized matrix is reduced to 9.2% of the original matrix’s range. Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Quantization Optimization (RQO). The RQO algorithm, as depicted in Figure 2, begins with the quantization of the original matrix to produce a quantized matrix. This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix. The residual matrix typically exhibits a significantly reduced data range, making it easier to compress. This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or QET.
@@ 3.4.1 Residual Quantization Optimization
3.4.2 Codebook Quantization Optimization
The codebook occupies a significant portion of the space in the QET algorithm. As shown in Table 1, our experiments indicate that when the compression ratio is 12, the codebook occupies 71.8% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space. Therefore, compressing the codebook can significantly reduce the space used by the algorithm. We proposed Codebook Quantization Optimization (CQO). Our optimization mainly employs the RTN method to compress the codebook.
@@ 4.1 Experiment setup
Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset. Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text. The results for the other datasets are provided in the appendix (Section LABEL:sec:appexponaddi). Below, we provide a detailed description of each dataset. Each element in the matrix is represented with 32 bits, except for the KV cache dataset, where each element is represented with 16 bits.
(1) Synthetic normal distribution dataset: This dataset was generated by drawing each element of the matrix from a truncated normal distribution with a mean of 0.5 and a standard deviation of 0.16. Additionally, one out of every ten thousand elements was replaced with an outlier value, randomly chosen between -100 and 100. For the synthetic normal distribution dataset, we generated three types of matrices of different sizes: synthetic dataset 1: 1024 $\times$ 128 matrices, synthetic dataset 2: 1024 $\times$ 512 matrices, and synthetic dataset 3: 1024 $\times$ 1024 matrices.
(2) LLM weight dataset: This dataset comprises weight matrices extracted from the large language model (LLM) LLaMA2 (Touvron et al. (2023)). The sizes of the LLM weight matrices are 11008 $\times$ 4096.
(3) KV cache dataset: The KV cache dataset is derived from the key-value pairs stored in the cache during inference in KV Quant (Hooper et al. (2024)). The key and value are stored in matrices, respectively. In the KV cache dataset, the sizes of both the K matrices and the V matrices are 4096 $\times$ 4096.
Platform and implementation: We conducted our algorithm evaluations on a high-performance server equipped with an Intel Xeon Platinum 6462C (Sapphire Rapids) processor, featuring 16 virtual CPUs (vCPUs), operating at a base frequency of 3.3 GHz, with a maximum turbo frequency of 3.9 GHz. The server also includes 64GB of memory, providing robust computational capabilities. All algorithms were implemented and executed on this server environment to ensure optimal performance for our experimental evaluations.
Metrics: We primarily measure the accuracy and time consumption of the algorithm. We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) as accuracy metrics. Quantization time ratio and Dequantization time ratio are used as time consumption metrics. Below, we introduce these metrics in detail.
@@
For time efficiency, we use Quantization Time (QT) and DeQuantization Time (DQT) as the metrics.
@@
Parameter Selection: We first introduce the parameter settings of the QET algorithm. For an $n\times d$ matrix with $a$ bit per element, we perform a quantization operation with a compression ratio of $\theta$. The QET algorithm iterates $l$ rounds. For residual optimization, we perform $r$ residual iterations. During the $i$th residual step, the number of clusters for centroids is $k_{i}$, and the number of groups during the grouping operation is $m$. For codebook quantization optimization, the bit length of the codebook after compression is $a^{\prime}$ bit. $R_{i}$ denotes the ratio of the space occupied by the codebook and the quantized matrix to the total space excluding the indicator maps during the $i$th iteration of residual optimization. To ensure that the total memory usage remains within the available memory limits, we have the following constraints on the parameters of the QET algorithm.
$k_{i}\times d\times a^{\prime}+2a+n\times m\times\log_{2}(k_{i})\leq\left[\frac{n\times d\times a}{\theta}-0.5\times n\times d\times l\right]\times R_{i},\quad\text{for }i=1,\ldots,r.$
Given the parameters $d$, $a^{\prime}$, $a$, $n$, $m$, $\theta$, $l$, and $R_{i}$, we calculate the value of $k_{i}$ using the bisection method. Unless otherwise specified, the QET parameters are configured as follows: $n$, $d$, and $a$ are determined by the input matrix. We perform a standard QET and one residual QET, both with codebook quantization. $r=2$, the first QET has 3 iterations, and the second QET does not perform matrix permutation ($l=3$). Additionally, $d/m=8$, $a^{\prime}=10$, $R_{1}=70\%$, and $R_{2}=30\%$.
For RTN, PQ, OPQ, and LOPQ, under the same space constraints, we use the same common parameters as our algorithm, while other parameters are set according to the recommended configurations from their respective papers. We define the names of different versions of the algorithm as follows: Vanilla refers to the basic QET algorithm without any optimizations. Vanilla_CQO refers to the Vanilla algorithm with codebook quantization optimization, where the codebook is quantized using the RTN algorithm. Vanilla_RQO indicates the Vanilla algorithm with residual optimization, where the residuals are processed using the QET algorithm. Finally, QET is the Vanilla algorithm with both residual optimization and codebook quantization optimization (including all optimizations), where the codebook is quantized using the RTN algorithm and the residuals are processed using the QET algorithm.
4.2 Matrix quantization result
In this section, we provide a comprehensive comparison of the accuracy and computational efficiency of the proposed QET algorithm against other benchmark methods. We will first discuss the precision metrics (MAE and MSE) across different datasets and compression ratios, followed by an analysis of the time metrics (QT and DQT) to evaluate the efficiency of each algorithm. The goal is to demonstrate the effectiveness of the QET algorithm in balancing accuracy with computational speed across various real and synthetic datasets. Some algorithms are unable to run on certain datasets or at specific compression ratios. For instance, in the case of square matrices, the rotation matrices for OPQ and LOPQ become as large as the original matrix, making compression impossible. In such cases, we observe missing data points or curves in the corresponding figures where the algorithms could not produce results.
(a) MAE of Synthetic 1.
(b) MSE of Synthetic 1.
(c) MAE of Synthetic 2.
(d) MSE of Synthetic 2.
(e) MAE of Synthetic 3.
(f) MSE of Synthetic 3.
(g) MAE of LLM.
(h) MSE of LLM.
(i) MAE of K cache.
(j) MSE of K cache.
(k) MAE of V cache.
(l) MSE of V cache.
Figure 3: MAE and MSE of different datasets.
MAE and MSE In our experiments, we evaluated the performance of our quantization algorithm across multiple datasets using MAE and MSE. These metrics provide a comprehensive understanding of the algorithm’s accuracy at different compression ratios, ranging from 2 to 16. The average MSE of RTN for the synthetic dataset 1, synthetic dataset 2, synthetic dataset 3, LLM dataset, K cache dataset, and V cache dataset is 71.74 times, 334.59 times, 247.39 times, 12.56 times, 10.57 times, and 7.69 times that of QET, respectively. Therefore, we do not include RTN in the figures for comparison.
For synthetic dataset 1: As shown in Figures  3(a) and  3(b) , our QET algorithm consistently outperforms the baseline algorithms across all compression ratios. The MAE and MSE values increase as the compression ratio rises, but QET maintains a significantly lower error compared to PQ and OPQ algorithms. The Vanilla version, a faster variant of QET, shows slightly higher error than the full QET version, but still remains competitive. For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 6.94%, while the Vanilla version reduces the MSE to 36.53%.
For Synthetic Dataset 2: In Figures  3(c) and  3(d), a similar trend is observed with Synthetic Dataset 2. The QET algorithm achieves the lowest MAE and MSE values, demonstrating its robustness across different types of synthetic data. Notably, the gap between QET and the other methods becomes more pronounced as the compression ratio increases, indicating the superiority of QET in handling higher compression rates without significantly sacrificing accuracy. Due to the space consumption of OPQ and LOPQ, they were unable to handle the compression of square matrices in this dataset. The vanilla version performs similarly to the QET algorithm. For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 7.13%, while the Vanilla version reduces the MSE to 36.51%.
For Synthetic Dataset 3: Figures  3(e) and  3(f) show the results for the third synthetic dataset. While the errors increase with higher compression ratios, QET still outperforms the other algorithms. The difference in performance becomes especially significant at higher compression ratios (e.g., 12 to 16), where QET and vanilla continue to show lower error rates. For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 7.21%, while the Vanilla version reduces the MSE to 36.54%.
For the LLM dataset, as depicted in Figures  3(g) and  3(h), the QET algorithm demonstrates superior performance. The MAE and MSE values for QET are consistently lower than those for Vanilla, PQ, and OPQ, indicating that QET is well-suited for compressing large-scale language model weights without significantly sacrificing accuracy. At higher compression ratios (13 to 16), the MAE and MSE of Vanilla are similar to those of QET. For example, at a compression ratio of 4, compared to the best-performing algorithm, PQ, QET reduces the MSE to 5.05%, while the Vanilla version reduces the MSE to 36.64%.
For KV Cache Dataset: The results of K cache is shown in Figures  3(i) and  3(j). The results of V cache is shown in Figures  3(k) and  3(l). Similar to the LLM datasets, the QET algorithm achieves the best results, with lower MAE and MSE values across all compression ratios. At compression ratios between 13 and 16, the MAE and MSE values of Vanilla closely approach those of QET. For example, at a compression ratio of 4 for K cache, compared to the best-performing algorithm, PQ, QET reduces the MSE to 13.33%, while the Vanilla version reduces the MSE to 50.24%. And at a compression ratio of 4 for V cache, compared to the best-performing algorithm, PQ, QET reduces the MSE to 11.89%, while the Vanilla version reduces the MSE to 45.83%.
In summary, the QET algorithm consistently outperforms baseline methods, achieving lower MAE and MSE values across all datasets and compression ratios. Vanilla also demonstrates relatively low MAE and MSE values, and on certain datasets, it performs comparably to QET at higher compression ratios.
(a) QT of Synthetic 1.
(b) DQT of Synthetic 1.
(c) QT of Synthetic 2.
(d) DQT of Synthetic 2.
(e) QT of Synthetic 3.
(f) DQT of Synthetic 3.
(g) QT of LLM.
(h) DQT of LLM.
(i) QT of K cache.
(j) DQT of K cache.
(k) QT of V cache.
(l) DQT of V cache.
Figure 4: QT (s) and DQT (s) of different datasets.
QT and DQT In this experiments, we evaluated the performance of the quantization process by measuring both Quantization Time (QT) and Dequantization Time (DQT) across multiple datasets, as shown in Figures 4. These metrics provide a comprehensive understanding of the computational efficiency of the algorithm at different compression ratios, ranging from 2 to 16.
For QT, the QET algorithm exhibits longer quantization times at lower compression ratios, but as the compression ratio increases, QT decreases significantly. This phenomenon is expected, as fewer centroids simplify the clustering process, resulting in faster compression as the compression ratio increases. Vanilla consistently achieves lower QT across all datasets compared to QET, making it the faster option, albeit with a slight loss in accuracy. In contrast, PQ and other baseline methods demonstrate moderate QT performance, although OPQ and LOPQ show slower performance at larger matrix sizes due to their complexity.
As for DQT, QET maintains a relatively stable performance across all compression ratios and datasets, with marginal increases in time as the compression ratio grows. Vanilla follows a similar trend but achieves faster dequantization. PQ remains competitive in terms of both QT and DQT.
4.3 Ablation Studies and Discussions
These experiments demonstrate the effectiveness of the algorithm optimizations. Specifically, both Residual Quantization Optimization (RQO) and Codebook Quantization Optimization (CQO) improve the baseline Vanilla algorithm and provide comparisons with the QET algorithm. The ablation studies primarily focus on accuracy improvements. We evaluated the performance on the LLM dataset by measuring both MAE and MSE.
(a) MAE of RQO.
(b) MSE of RQO.
(c) MAE of CQO.
(d) MSE of CQO.
Figure 5: Ablation Studies of RQO and CQO.
Residual Quantization Optimization (RQO) As illustrated in Figures 5(b) and LABEL:fig:ablationrqomse, the introduction of RQO into the Vanilla algorithm leads to a notable reduction in MAE and MSE values. At lower compression ratios, Vanilla with RQO achieves performance that is very close to QET, particularly in terms of MSE. The gap widens at higher compression ratios because, while the storage space remains the same for both algorithms, RQO requires more codebooks due to the multi-layer structure, which reduces the number of centroids in clustering. On the other hand, QET optimizes codebook compression, allowing for more centroids and better MAE and MSE performance at higher compression ratios. For instance, at a compression ratio of 4, the RQO optimization reduces MSE to 16.23% of the original value.
Codebook Quantization Optimization (CQO)
In Figures 5(c) and 5(d), the impact of CQO on the Vanilla algorithm is shown. CQO helps to decrease MAE and MSE, especially at lower compression ratios. CQO optimization effectively reduces the size of the codebook, leading to improvements in both MAE and MSE. However, at higher compression ratios, the accuracy improvements become less noticeable because the codebook size is already small at these higher compression levels, leaving limited room for further compression. For instance, at a compression ratio of 4, the CQO optimization reduces MSE to 40.08% of the original value.